# Linear analysis of I-C-Love universal relations for neutron stars

### v1

@@ Abstract
Neutron stars (NSs) are excellent laboratories for testing gravity theories as they are strongly self-gravitating bodies and have rich observational phenomena. However, strong-gravity effects in NS could be degenerate with their equation of state (EOS) which is largely unknown. Fortunately, there exist the so-called universal relations among the NS macroscopic quantities that are found to be insensitive to the underlying EOS. Studying the origin of these relations can lead to a better understanding of NSs and the gravitational interaction. We develop a new perspective of view to analyze the I-C and I-Love universal relations for NSs. At the linear order, we separate the deviation of the universal relations into two parts, where one is the EOS perturbation while the other only depends on the background star structures. The smallness of the second part fully determines the universality of the relation irrespective of the form of the first part. We discuss the validity of our linear approximation when considering the difference among realistic EOSs. Our study can be regarded as a new frame for quantitative representation of the universality and may provide new insights to the universal relations of NSs.
@@ I Introduction
Neutron stars (NSs) are extremely compact objects in the Nature. Thanks to their strong self-gravity, large density in the inner region, and abundant observational phenomena, they are ideal laboratories for studying fundamental physics, including the strong-field gravity and nuclear physics Lattimer and Prakash (2001); Shao (2023); Freire and Wex (2024). For example, the mass-radius relationship of NSs strongly depends on the supranuclear matter equation of state (EOS), which is the relation between the matter density $\rho$ and pressure $p$. Thus, independently measuring the mass $M$ and radius $R$ of a NS can provide valuable constraints on the EOS at the high density regime that is unachievable with terrestrial nuclear experiments. The observations of heavy pulsars with masses $\gtrsim 2\,M_{\odot}$ have already ruled out those EOSs that are too “soft” for supporting a large mass NS in the general relativity (GR) Fonseca et al. (2021); Demorest et al. (2010); Antoniadis et al. (2013). Current X-ray observations with the NICER satellite can measure the mass and radius of a pulsar to $\sim 10\%$ accuracy and have also been used to constrain the EOS Riley et al. (2019); Raaijmakers et al. (2019).
@@
The degeneracy between the strong-field gravity effects and the nuclear matter EOS in NSs causes a challenge for studying both aspects Shao (2019). A test of strong-field gravity can only be done if we assume that we know the underlying EOS or systematically taken into account its large uncertainties, and vise versa. One way to break this degeneracy is provided by the finding of the so-called universal relations among some NS properties Yagi and Yunes (2013a, b); Pappas and Apostolatos (2014); Chakrabarti et al. (2014); Gao et al. (2023); Luk and Lin (2018); Shao and Yagi (2022). These relations, for example, the well-known I-Love-Q relation for the dimensionless NS moment of inertia $\tilde{I}$, tidal Love number $\tilde{\Lambda}$ and spin induced quadrupole moment $\tilde{Q}$, are largely insensitive to the EOSs but sometimes do depend on the gravity theories or even not universal in theories beyond GR Yagi and Yunes (2017). Thus, independent measurement of these quantities will enable us to have EOS independent test of gravity theories. In this regard, the I-Love relation will be especially interesting as the tidal Love number of NS can be measured with gravitational wave observations Abbott et al. (2018), while it is expected to measure the NS moment of inertia with the timing observation of the Double Pulsar J0737--3039A/B Hu et al. (2020); Kramer et al. (2021).
Besides the interesting applications of these universal relations, studying themselves also provides us better understanding of NSs and gravity theories. The origin of the universality, which can be as good as $\lesssim 1\%$ level in the I-Love-Q relations, is still in discussion. In literature, it is proposed at the beginning that the universality may come from the similarity of the different EOSs in the low density regime or it is the remnants of the no-hair theorem Yagi and Yunes (2013b, a). Considering that some relations like I-Love-Q even hold among NSs and quark stars (QSs), which are composed with quark matter that has a completely different EOS behavior in the low density regime, it is then proposed that the universality may come from the emergent self-similarity of isodensity contours inside NSs Yagi et al. (2014) or the realistic EOS is “close” to the incompressible limit Sham et al. (2015); Katagiri et al. (2025). There are also analytical calculations of these relations based on series expansion of the star’s compactness and assuming an incompressible EOS Chan et al. (2015) or Tolman VII model that phenomenologically describes the density profile inside NSs Jiang and Yagi (2020); Lowrey et al. (2025).
@@
In this work, we analyze the NSs’ I-C and I-Love universal relations from a new perspective of view. Rather than parametrizing or expanding the EOSs or NS structures, we consider the linear response of the universal relation under any EOS perturbations. With this approach, we are able to separate the deviation of the universal relation caused by EOS perturbations into two parts: one is the difference in the EOSs and the other factor only depends on the background star solution. Thus, the smallness of the second factor fully accounts for the universality. Our approach also does not assume an expansion in the star’s compactness thus is applicable for both Newtonian and GR configurations.
@@ II I-C universal relation
In Fig. 1 we show the I-C universal relation for several realistic EOSs including NSs and QSs as well as polytropic EOSs with $\rho(p)=\rho_{n}(p/\rho_{n})^{n/(n+1)}$. The dimensionless moment of inertial is defined to be $\tilde{I}\equiv I/M^{3}$ and the compactness is $C=M/R$ Yagi and Yunes (2013a, b). As we focus on the dimensionless relation, the dimensionful quantity $\rho_{n}$ dose not affect the final relation and we choose it to be $1$ as the unit. From the figure one can see that for realistic EOSs, the relation holds well to the level of several percentages. However, this relation has different branches for NSs and QSs. As expected, the polytropic EOS with $n=1$ is a good approximation for realistic NS EOSs while $n=0$ can describe the EOSs of QSs.
@@ III Linear response of the I-C relation
and thus $R$, $M$, $I$ for the star at $p=0$. With the background solution, considering another EOS given by
@@
One can then obtain the differential equations for $\delta r$, $\delta m$ and $\delta i$ by substituting the above equations to the NS structure equations. Keeping only the leading order terms in the small quantities, one has
@@
Here we use “$\Delta$” to distinguish it from the perturbation of EOS. To deal with this perturbation, we can integrate one step outwards to get
@@ IV Analytical solution to a constant-density star in the Newtonian gravity
It is suggest that the universal relations may come from the similarity of the realistic EOSs and an incompressible, constant-density EOS Sham et al. (2015). Previous study also showed that, the I-Love relation is independent of leading order EOS perturbation around the incompressible EOS in the Newtonian limit (at the relative leading order in $C$Chan et al. (2016). Thus here we study the I-C relation for a constant-density star in the Newtonian gravity analytically as an illustration.
@@ V Linear response for compact stars with realistic EOSs
In this section, we focus on NSs and QSs with realistic EOSs in GR. To obtain the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ for realistic EOSs, in general one needs to numerically solve Eq. (17). In Fig. 3, we show this factor for NSs with the AP4 EOS Akmal et al. (1998) as well as QS with the SQM1 EOS Prakash et al. (1995). As in general this factor depends on the background stars, we plot the NS cases for three different star compactnesses. We also plot Eq. (43) for comparison. One can see that, the Newtonian constant density approximation does capture the order of magnitude for the factor in realistic cases, except for a small region where $p_{1}/p_{0}$ is close to zero, that is, the outer region of the NS. We find that when $p_{1}/p_{0}$ tends to zero, this factor grows up towards infinity (or towards a relatively large but finite number in the case of QSs) for realistic EOSs. We will discuss this behavior later in detail as it is closely related to the fact that NSs and QSs have different branches of the I-C relation.
@@
To understand this divergent behavior, we notice that in Eq. (40), the behavior of the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ depends on $\delta I(p_{1})/I$, $\delta M(p_{1})/M$, and $\delta R(p_{1})/R$, which are directly solved from the perturbation equations. In Fig. 7, we plot these factors separately. From this figure, one can clear see that the divergent behavior at the surface of the star is dominated by the factor $\delta R(p_{1})/R$. This divergent behavior can then be seen from the initial condition. Simply taking the Newtonian approximated condition (33), one can see that, close to the surface of the star, $\delta r/r\sim-1/C_{1}\rho_{1}^{2}$, where $C_{1}\equiv m_{1}/r_{1}$ can be approximated to the star’s compactness $C$. $\delta R(p_{1})/R$ evolves from the initial condition at $p_{1}$ to the surface of the star. However, when $p_{1}$ is close to zero, the evolution does not affect the value significantly. Thus, for NSs, the factor $1/\rho_{1}^{2}$ and $\delta R(p_{1})/R$ are divergent when $p_{1}/p_{0}$ goes to zero. In Fig. 7 we also plot this initial condition with an approximation that $C_{1}\sim 0.2$. It can be seen that this Newtonian initial condition already captures the behavior for small $p_{1}/p_{0}$. Differently for QSs, also for the constant-density star we considered before, they have a finite density at the surface of the star. All the factors remain finite there. In fact, more detailed behaviors, like that $\delta I(p_{1})/I$ and $\delta M(p_{1})/M$ go to zero, as shown in Fig. 8, can also be read from the initial condition.
@@
Finally, the integrated factor $p_{0}\rho_{0}{\rm D}\tilde{I}/\tilde{I}$ is divergent if one considers a NS background but takes a QS EOS as the perturbed EOS. As the integration at small $p_{1}$ is roughly $\int 1/\rho_{1}^{2}{\rm d}p_{1}$, for a NS EOS, the low density part can be approximated by $\rho\propto p^{5/3}$ Read et al. (2009) and thus leads to a divergent integral. As we will see later, this divergence means that the linear approximation fails for this situation, suggesting an intrinsic difference in the I-C relation for NSs and QSs. We also note that, the equivalence between taking the NS as background and taking QS as background only holds when both of them give finite results.
@@ VII Validity of the linear order analysis
will provide exactly the same relation as the EOS $\rho(p)$, where $\alpha$ is any positive constant. We have already encountered this fact for polytropic EOSs as we choose $\rho_{n}=1$. This scaling freedom means that a direct comparison between stars with different EOSs but the same central pressure as we did before might not be the proper choice. A better choice is to compare the star with, for example, the same $C$ or $\tilde{\Lambda}$ depending on the relation we want to study. At the linear order, denoting $\alpha=1+\Delta$ where $|\Delta|\ll 1$, we have
$\rho^{\prime}(p)=\rho(p)+\left(\rho-\frac{{\rm d}\rho}{{\rm d}p}p\right)\Delta\,.$ (62)
Any proper comparison at the linear order should modulo such a scaling effect. Nevertheless, as we argued above, the scaling leaves the universal relations unchanged, thus the quantity defined in Eq. (21) or Eq. (49) should be zero for EOS perturbations have the above form by definition. As an example, one can directly verify that the integral (43) is zero as the argument also holds in the Newtonian gravity. This independence means that the factor we use is in fact a proper factor for linear-order analysis.
For comparing realistic EOSs, we suggest that a proper scaling needs to be done before applying the linear analysis. From previous figures we find that, the linear prediction of $\delta\tilde{I}$ is in fact pretty accurate even for realistic EOS differences. Thus, a scaling that leads to the comparison of two stars with a similar $C$ or $\tilde{\Lambda}$ should be well described by the linear response.

### v2

@@ Abstract
Neutron stars (NSs) are excellent laboratories for testing gravity theories as they are strongly self-gravitating bodies and have rich observational phenomena. However, strong-field gravity effects in NS could be degenerate with their equation of state (EOS) which is largely unknown. Fortunately, there exist the so-called universal relations among the NS macroscopic quantities that are found to be insensitive to the underlying EOS. Studying the origin of these relations can lead to a better understanding of NSs and the gravitational interaction. We develop a new perspective of view to analyze the I-C and I-Love universal relations for NSs. At the linear order, we separate the deviation of the universal relations into two factors that are multiplied together. One is the EOS difference while the other factor only depends on the background star structure. The smallness of the second factor that does not depend on the EOS difference then indicates the origin of the universality. We discuss the validity of our linear approximation when considering the difference among realistic EOSs. Our study can be regarded as a new frame for quantitative representation of the universality and may provide new insights to the universal relations of NSs.
@@ I Introduction
Neutron stars (NSs) are extremely compact objects in the Nature. Thanks to their strong self-gravity, large density in the inner region, and abundant observational phenomena, they are ideal laboratories for studying fundamental physics, including the strong-field gravity and nuclear physics Lattimer and Prakash (2001); Shao (2023); Freire and Wex (2024). For example, the mass-radius relationship of NSs strongly depends on the supranuclear matter equation of state (EOS), which is the relation between the matter density $\rho$ and pressure $p$. Thus, independently measuring the mass $M$ and radius $R$ of a NS can provide valuable constraints on the EOS at the high density regime that is unachievable with terrestrial nuclear experiments. The observations of heavy pulsars with masses $\gtrsim 2\,M_{\odot}$ have already ruled out those EOSs that are too “soft” for supporting a large mass NS in the general relativity (GR) Fonseca and others (2021); Demorest et al. (2010); Antoniadis and others (2013). Current X-ray observations with the NICER satellite can measure the mass and radius of a pulsar to $\sim 10\%$ accuracy and have also been used to constrain the EOS Riley and others (2019); Raaijmakers and others (2019).
@@
The degeneracy between the strong-field gravity effects and the nuclear matter EOS in NSs causes a challenge for studying both aspects Shao (2019). A test of strong-field gravity can only be done if we assume that we know the underlying EOS or systematically taken into account its large uncertainties, and vise versa. One way to break this degeneracy is provided by the finding of the so-called universal relations among some NS properties Yagi and Yunes (2013b, a); Pappas and Apostolatos (2014); Chakrabarti et al. (2014); Gao et al. (2023); Luk and Lin (2018); Shao and Yagi (2022). These relations, for example, the well-known I-Love-Q relation for the dimensionless NS moment of inertia $\tilde{I}$, tidal Love number $\tilde{\Lambda}$ and spin induced quadrupole moment $\tilde{Q}$, are largely insensitive to the EOSs but sometimes do depend on the gravity theories or even not universal in theories beyond GR Yagi and Yunes (2017). Thus, independent measurement of these quantities will enable us to have EOS independent test of gravity theories. In this regard, the I-Love relation will be especially interesting as the tidal Love number of NS can be measured with gravitational wave observations Abbott and others (2018), while it is expected to measure the NS moment of inertia with the timing observation of the Double Pulsar J0737-3039A/B Hu et al. (2020); Kramer and others (2021).
Besides the interesting applications of these universal relations, studying themselves also provides us better understanding of NSs and gravity theories. The origin of the universality, which can be as good as $\lesssim 1\%$ level in the I-Love-Q relations, is still in discussion. In literature, it is proposed at the beginning that the universality may come from the similarity of the different EOSs in the low density regime or it is the remnants of the no-hair theorem Yagi and Yunes (2013a, b). Considering that some relations like I-Love-Q even hold among NSs and quark stars (QSs), which are composed with quark matter that has a completely different EOS behavior in the low density regime, it is then proposed that the universality may come from the emergent self-similarity of isodensity contours inside NSs Yagi et al. (2014) or the realistic EOS is “close” to the incompressible limit Sham et al. (2015); Katagiri et al. (2025). There are also analytical calculations of these relations based on series expansion of the star’s compactness and assuming an incompressible EOS Chan et al. (2015) or Tolman VII model that phenomenologically describes the density profile inside NSs Jiang and Yagi (2020); Lowrey et al. (2025).
@@
In this work, we analyze the NSs’ I-C and I-Love universal relations from a new perspective of view. Rather than parametrizing or expanding the EOSs or NS structures, we consider the linear response of the universal relation under any EOS perturbations. With this approach, we are able to separate the deviation of the universal relation caused by EOS perturbations into two parts that are multiplied together: one is the difference in the EOSs and the other factor only depends on the background star solution. As one can have almost arbitrary EOS difference if the universal relation holds, the smallness of the second factor fully accounts for the universality. Our approach also does not assume an expansion in the star’s compactness thus is applicable for both Newtonian and GR configurations.
@@ II I-C universal relation
In Fig. 1 we show the I-C universal relation for several realistic EOSs including NSs and QSs as well as polytropic EOSs with $\rho(p)=\rho_{n}(p/\rho_{n})^{n/(n+1)}$. The dimensionless moment of inertial is defined to be $\tilde{I}\equiv I/M^{3}$ and the compactness is $C=M/R$ Yagi and Yunes (2013b, a). As we focus on the dimensionless relation, the dimensionful quantity $\rho_{n}$ dose not affect the final relation and we choose it to be $1$ as the unit. From the figure one can see that for realistic EOSs, the relation holds well to the level of several percentages. However, this relation has different branches for NSs and QSs. As expected, the polytropic EOS with $n=1$ is a good approximation for realistic NS EOSs while $n=0$ can describe the EOSs of QSs.
@@ III Linear response of the I-C relation
and thus $R$, $M$, II for the star at $p=0$. With the background solution, considering another EOS given by
@@
Assuming $\delta\rho(p)$ is an EOS perturbation satisfying $\delta\rho(p)/\rho(p)\ll 1$ for $p\in[0,p_{0}]$, one can then obtain the linearized differential equations for $\delta r$, $\delta m$ and $\delta i$ by substituting the above equations to the NS structure equations and only keeping the leading order terms. The linearization also requires the variable set $\delta r(p)/r(p)$, $\delta m(p)/m(p)$ and $\delta i(p)/i(p)$ to be small, which is in general guaranteed by the assumption $\delta\rho(p)/\rho(p)\ll 1$. Note that at $p=p_{0}$ the condition $r=m=i=0$ is only a coordinate singularity. At the center of the star, one can solve the TOV equations to obtain
$r(p)=\sqrt{\frac{3}{2\pi}\frac{p_{0}-p}{(\rho_{0}+p_{0})(\rho_{0}+3p_{0})}}+O(p_{0}-p)^{3/2}\,$, (13)
where $\rho_{0}=\rho(p_{0})$. Therefore,
$\left.\frac{\delta r}{r}\right|_{p=p_{0}}=-\frac{\rho_{0}+2p_{0}}{(\rho_{0}+p_{0})(\rho_{0}+3p_{0})}\delta\rho(p_{0})\,$, (14)
which is also small if $\delta\rho(p_{0})/\rho(p_{0})\ll 1$; it is similar for $\delta m/m$ and $\delta i/i$.
The linearized equations for the $\delta$-quantities read
@@
Before further proceeding, here we note that the dimension of the delta function source $\delta\rho=\delta(p-p_{1})$ is $p^{-1}$ instead of $\rho$ according to the definition of delta function. Therefore, the corresponding solution like $\delta r$ has a dimension of $r/p\rho$ instead of $r$, which might be less intuitive. Nevertheless, physical quantities that are obtained by applying superposition principle have the correct dimension as shown below.
@@
Here we use $\Delta$” to distinguish it from the perturbation of EOS. To deal with this perturbation, we can integrate one step outwards to get
@@ IV Analytical solution to a constant-density star in the Newtonian gravity
It is suggest that the universal relations may come from the similarity of the realistic EOSs and an incompressible, constant-density EOS Sham et al. (2015). Previous study also showed that, the I-Love relation is independent of leading order EOS perturbation around the incompressible EOS in the Newtonian limit (at the relative leading order in CCChan et al. (2016). Thus here we study the I-C relation for a constant-density star in the Newtonian gravity analytically as an illustration.
@@ V Linear response for compact stars with realistic EOSs
In this section, we focus on NSs and QSs with realistic EOSs in GR. We use the EOS data provided by Lattimer and Prakash (2001), which are in a tabulated form. The EOSs for NS have a crust EOS down to a density of $10^{8}\,{\rm g\,cm^{-3}}$. To obtain the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ for realistic EOSs, in general one needs to numerically solve Eq. (19). In Fig. 3, we show this factor for NSs with the AP4 EOS Akmal et al. (1998) as well as QS with the SQM1 EOS Prakash et al. (1995). As in general this factor depends on the background stars, we plot the NS cases for three different star compactnesses. We also plot Eq. (45) for comparison. One can see that, the Newtonian constant density approximation does capture the order of magnitude for the factor in realistic cases, except for a small region where $p_{1}/p_{0}$ is close to zero, that is, the outer region of the NS. We find that when $p_{1}/p_{0}$ tends to zero, this factor grows up towards infinity (or towards a relatively large but finite number in the case of QSs) for realistic EOSs. We will discuss this behavior later in detail as it is closely related to the fact that NSs and QSs have different branches of the I-C relation.
@@
To understand this divergent behavior, we notice that in Eq. (42), the behavior of the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ depends on $\delta I(p_{1})/I$, $\delta M(p_{1})/M$, and $\delta R(p_{1})/R$, which are directly solved from the perturbation equations. In Fig. 7, we plot these factors separately. From this figure, one can clear see that the divergent behavior at the surface of the star is dominated by the factor $\delta R(p_{1})/R$.
The divergence of $\delta R(p_{1})$ near the surface of the NS can be analytically estimated via a polytropic crust EOS model. Considering EOS perturbations in the low pressure region, the inner structure of the star then is irrelevant to our problem and the effect can be fully characterized by the statement that, at $p=p_{R}$ and $\rho=\rho_{R}$, the core of the star has mass $M$, radius $R$, where $p_{R}$ is some small pressure. We can then assume that for $p<p_{R}$, the NS crust has a polytropic EOS that $\rho(p)=\rho_{R}(p/p_{R})^{n/(n+1)}$ and we can model the crust of the NS with Newtonian approximation,
$\frac{{\rm d}r}{{\rm d}p}=-\frac{1}{\rho}\frac{r^{2}}{m}\,.$ (46)
Further, it is known that the crust of NS contains only a small amount of mass but provides a relatively considerable radius Chan et al. (2015), therefore, when considering the behavior of $\delta r$, we can set $m=M$ in the above equation (also note that $\delta r$ is the only divergent part so we can safely ignore $\delta m$ there).
With these approximation, one can solve the structure of the curst analytically, which reads
$\frac{1}{r}-\frac{1}{R}=(n+1)\frac{p_{R}^{\frac{n}{n+1}}}{M\rho_{R}}\left(p^{\frac{1}{n+1}}-p_{R}^{\frac{1}{n+1}}\right)\,.$ (47)
This gives a crust with a depth $z\approx(n+1)R^{2}p_{R}/M\rho_{R}$. Following the procedure introduced in Sec. III, one can then write down the equation for $\delta r(p)$ caused by an EOS perturbation $\delta\rho(p)=\delta(p-p_{1})$ as
$\frac{{\rm d}\delta r}{{\rm d}p}=-\frac{2r}{\rho M}\delta r\,$, (48)
with initial condition $\delta r(p_{1})=-r(p_{1})^{2}/\rho(p_{1})^{2}M$. This equation again can be integrated analytically and one can obtain
$\delta R(p_{1})=\delta r(p_{1})\left(\frac{\frac{M\rho_{R}/R-p_{R}}{p_{R}^{n/(n+1)}}+p_{1}^{1/(n+1)}}{\frac{M\rho_{R}/R-p_{R}}{p_{R}^{n/(n+1)}}}\right)^{2}\,\approx-\frac{R^{2}}{M\rho(p_{1})^{2}}.$ (49)
Thus, near the surface $\delta R(p_{1})/R$ diverges as $\rho(p_{1})^{-2}$. In Fig. 7, we also plot the above estimation with $M/R\approx 0.2$. It can be seen that this estimation captures the divergent behavior of $\delta R(p_{1})/R$. Differently for QSs, near the surface, they behave more like the constant density star as we discussed before. Factors remain finite there.
We should note that, the divergence of $\delta R(p_{1})$ is a physical result and itself does not cause problem to our linear analysis. The quantity that needs to be finite and small is the integral defined in Eq. (22) when the linear approximation holds. As an example, we may consider the EOS perturbation with $n\rightarrow n+\epsilon$ in the above crust model. The corresponding change in the density is
$\delta\rho(p)=\frac{\epsilon\rho_{R}}{(n+1)^{2}}\left(\frac{p}{p_{R}}\right)^{\frac{n}{n+1}}\ln\frac{p}{p_{R}}\,$, (50)
which is well defined for $n>0$. The integration,
$\int_{0}^{p_{R}}{\rm d}p_{1}\delta R(p_{1})\delta\rho(p_{1})\approx\epsilon\frac{R^{2}p_{R}}{M\rho_{R}}\,$, (51)
is just the change in $z$ that $\delta z=\delta nR^{2}p_{R}/M\rho_{R}=\epsilon R^{2}p_{R}/M\rho_{R}$. One can also notice that, in Fig. 6, the integration of ${\rm D}\tilde{I}$ is finite for NS-NS comparison. As $\delta R(p_{1})$ is the only divergent part in ${\rm D}\tilde{I}$, the integration of $\delta R(p_{1})$ for realistic NS EOSs, remains finite.
@@
Finally, the integrated factor $p_{0}\rho_{0}{\rm D}\tilde{I}/\tilde{I}$ is divergent if one considers a NS background but takes a QS EOS as the perturbed EOS. As the QS EOS has a finite surface density $\rho_{s}$ while NS EOS at the low density part can be approximated by $\rho\propto p^{3/5}$, the integration at small $p_{1}$ is roughly $\int 1/\rho_{1}^{2}{\rm d}p_{1}$ and divergent. As we will see in Sec. VII, this divergence means that the linear approximation fails for this situation, suggesting an intrinsic difference in the I-C relation for NSs and QSs. We also note that, the equivalence between taking any one of the star as background only holds when both choices give finite results and it does not apply to the NS-QS comparison.
@@ VII Validity of the linear order analysis
will provide exactly the same relation as the EOS $\rho(p)$, where $\alpha$ is any positive constant. We have already encountered this fact for polytropic EOSs as we choose $\rho_{n}=1$. Therefore, a proper scaling may be applied before the linear analysis as the scaling changes the meaning of the same central pressure. In fact, for I-Love relation, a comparison should be made between the dimensionless moment of inertia of two stars with different EOSs but the same dimensionless tidal deformability. The quantities ${\rm D}\tilde{I}$ we defined in Eq. (57), originating from the rule of partial differentiation
$\left.\frac{\partial\tilde{I}(\tilde{\Lambda}(p_{c},\mathrm{EOS}),\mathrm{EOS})}{\partial\mathrm{EOS}}\right|_{\tilde{\Lambda}}=\left.\frac{\partial\tilde{I}}{\partial{\mathrm{EOS}}}\right|_{p_{c}}-\left.\frac{\partial\tilde{I}}{\partial\tilde{\Lambda}}\right|_{\mathrm{EOS}}\left.\frac{\partial\tilde{\Lambda}}{\partial\mathrm{EOS}}\right|_{p_{c}}\,$, (70)
only provide such a comparison at the linear order. It is possible that realistic EOSs that largely deviate from each other might be similar after scaling. Therefore, for comparing realistic EOSs, we suggest that a proper scaling needs to be done before applying the linear analysis. From previous figures we find that, the linear prediction of $\delta\tilde{I}$ is in fact pretty accurate even for realistic EOS differences. Thus, a scaling that leads to the comparison of two stars with a similar $C$ or $\tilde{\Lambda}$ should be well described by the linear response.